### v1

# Intersections of sumsets in additive number theory

Abstract.
Let AA be a subset of an additive abelian semigroup and let hAhA be the $h$-fold sumset of AA. The following question is considered: Let $(A_{q})_{q=1}^{\infty}$ be a strictly decreasing sequence of sets in the semigroup and let $A=\bigcap_{q=1}^{\infty}A_{q}$. When does one have
$hA=\bigcap_{q=1}^{\infty}hA_{q}$
for some or all $h\geq 2$?
Key words and phrases:
Sumset, Haar measure of sumsets, additive number theory, combinatorial number theory
2000 Mathematics Subject Classification:
11B13, 11B05, 11B75, 11P70, 22D99
Supported in part by PSC-CUNY Research Award Program grant 66197-00 54.
1. The sumset intersection problem
Let $\mathbf{N}$, $\mathbf{N}_{0}$, and $\mathbf{Z}$ denote, respectively, the sets of positive integers, nonnegative integers, and integers.
A sequence of sets $(A_{q})_{q=1}^{\infty}$ is decreasing if $A_{q}\supseteq A_{q+1}$ for all $q\in\mathbf{N}$. A decreasing sequence of sets $(A_{q})_{q=1}^{\infty}$ is strictly decreasing if $A_{q}\neq A_{q+1}$ for all $q\in\mathbf{N}$ and asymptotically strictly decreasing if the sequence is not eventually constant, that is, if $A_{q}\neq A_{q+1}$ for infinitely many $q$.
Let AA be a subset of an additive abelian semigroup. The $h$-fold sumset of AA is the set
$hA=\{a_{1}+\cdots+a_{h}:a_{i}\in A\text{ for all }i=1,\ldots,h\}.$
Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of sets in the semigroup and let
$A=\bigcap_{q=1}^{\infty}A_{q}.$
For all $q\in\mathbf{N}$, we have $A\subseteq A_{q}$ and so $hA\subseteq hA_{q}$. It follows that
(1) $hA=h\left(\bigcap_{q=1}^{\infty}A_{q}\right)\subseteq\bigcap_{q=1}^{\infty}hA_{q}.$
This inclusion can be proper. For example, in the semigroup $\mathbf{Z}$ of integers, if $q\in\mathbf{N}$ and
$A^{\sharp}_{q}=\{r\in\mathbf{Z}:|r|\geq q\}$
then
$n=(n+q)+(h-2)q+(-(h-1)q)\in hA^{\prime}$
and
n=(nq)+(h2)(q)+((h1)q)-n=(-n-q)+(h-2)(-q)+((h-1)q)
for all $n\in\mathbf{N}_{0}$, and so $hA^{\sharp}_{q}=\mathbf{Z}$ for all $h\geq 2$. Let AA be any finite set of integers. The sumset hAhA is finite for all $h\in\mathbf{N}$. The asymptotically strictly decreasing sequence of sets $(A_{q})_{q=1}^{\infty}$ defined by
$A_{q}=A\cup A^{\sharp}_{q}=A\cup\{r\in\mathbf{Z}:|r|\geq q\}$
satisfies $A=\bigcap_{q=1}^{\infty}A_{q}$. We have $hA_{q}=\mathbf{Z}$ for all $q\in\mathbf{N}$ and so
$hA\neq\mathbf{Z}=\bigcap_{q=1}^{\infty}hA_{q}.$
We want to know when equality replaces inclusion in relation (1).
Problem 1.
Describe, classify, understand the strictly decreasing sequences $(A_{q})_{q=1}^{\infty}$ of subsets of an additive abelian semigroup such that
(2) $h\left(\bigcap_{q=1}^{\infty}A_{q}\right)=\bigcap_{q=1}^{\infty}hA_{q}.$
For related work, see [1][11].
2. Intersections of sets of integers
Theorem 1.
Let AA be a set of nonnegative integers. Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of sets of nonnegative integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$. For every positive integer $h$,
$hA=\bigcap_{q=1}^{\infty}hA_{q}.$
Proof.
We must show that if $n\in\bigcap_{q=1}^{\infty}hA_{q}$, then $n\in hA$.
The number of representations of $n$ as a sum of $h$ nonnegative integers is finite. Indeed, the set $\mathcal{R}(n)$ of $h$-tuples $(b_{1},\ldots,b_{h})\in\mathbf{N}_{0}^{h}$ such that
$n=b_{1}+\cdots+b_{h}$
satisfies $|\mathcal{R}(n)|=\binom{n+h-1}{h-1}$. For all $q\in\mathbf{N}$, we have $n\in hA_{q}$ and so there exists $(a_{1,q},\ldots,a_{h,q})\in A_{q}^{h}\cap\mathcal{R}(n)$ such that
$n=a_{1,q}+\cdots+a_{h,q}.$
Because $\mathcal{R}(n)$ is finite, there is some $h$-tuple $(b_{1},\ldots,b_{h})\in\mathcal{R}(n)$ such that
$(b_{1},\ldots,b_{h})=(a_{1,q},\ldots,a_{h,q})$
for infinitely many $q$ and so $b_{i}=a_{i,q}\in A_{q}$ for infinitely many $q$. Because the sequence $(A_{q})_{q=1}^{\infty}$ is decreasing, we have $b_{i}\in A_{q}$ for all $q$ and so $b_{i}\in\bigcap_{q=1}^{\infty}A_{q}=A$ and $n=b_{1}+\cdots+b_{h}\in hA$. This completes the proof. ∎
The set AA of integers is a basis of order $h$ for $\mathbf{Z}$ if $hA=\mathbf{Z}$. The set AA is a basis for $\mathbf{Z}$ if AA is a basis of order $h$ for $\mathbf{Z}$ for some positive integer $h$. If AA is a basis of order $h$ for $\mathbf{Z}$ and if $A\subseteq A^{\prime}$, then $A^{\prime}$ is a basis of order $h^{\prime}$ for $\mathbf{Z}$ for all $h^{\prime}\geq h$.
Theorem 2.
Let AA be a set of integers.
  1. (1)
    If AA is a basis of order $h_{0}$ for $\mathbf{Z}$ and if $(A_{q})_{q=1}^{\infty}$ is any decreasing sequence of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$, then
    $hA=\bigcap_{q=1}^{\infty}hA_{q}$
    for all $h\geq h_{0}$.
  2. (2)
    If AA is not a basis of order $h_{0}$ for $\mathbf{Z}$, then there exists a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ but
    $h_{0}A\neq\bigcap_{q=1}^{\infty}hA_{q}.$
  3. (3)
    If AA is not a basis for $\mathbf{Z}$, then there exists a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ but
    $hA\neq\bigcap_{q=1}^{\infty}hA_{q}$
    for all $h\in\mathbf{N}$.
Proof.
(1) Let AA be a basis of order $h_{0}$ for $\mathbf{Z}$ and let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$. For all $h\geq h_{0}$ and for all $q\in\mathbf{N}$, we have $hA=hA_{q}=\mathbf{Z}$ and so
$hA=\mathbf{Z}=\bigcap_{q=1}^{\infty}hA_{q}.$
(2) If AA is not a basis of order $h_{0}$ for $\mathbf{Z}$, then $h_{0}A\neq\mathbf{Z}$. For all $q\in\mathbf{N}$, let
$A_{q}=A\cup A^{\sharp}=A\cup\{r\in\mathbf{Z}:|r|\geq q\}.$
Then $A=\bigcap_{q=1}^{\infty}A_{q}$ and $h_{0}A_{q}=\mathbf{Z}$ for all $q\in\mathbf{N}$. It follows that
$h_{0}A\neq\mathbf{Z}=\bigcap_{q=1}^{\infty}h_{0}A_{q}.$
(3) If AA is not a basis for $\mathbf{Z}$, then $hA\neq\mathbf{Z}$ for all $h\in\mathbf{N}$. For all $q\in\mathbf{N}$, let
$A_{q}=A\cup A^{\sharp}=A\cup\{r\in\mathbf{Z}:|r|\geq q\}.$
Then $A=\bigcap_{q=1}^{\infty}A_{q}$ and $h_{0}A_{q}=\mathbf{Z}$ for all $q\in\mathbf{N}$. It follows that
$hA\neq\mathbf{Z}=\bigcap_{q=1}^{\infty}hA_{q}$
for all $h\in\mathbf{N}$. This completes the proof. ∎
Theorem 3.
Let AA be an infinite set of integers that is bounded below and that does not contain all sufficiently large integers. There exists an asymptotically strictly decreasing sequence $(A_{q})_{q=1}^{\infty}$ such that $hA=\bigcap_{q=1}^{\infty}hA_{q}$ for all $h\in\mathbf{N}$.
Proof.
By Theorem 1, we can assume that $a^{*}=\min(A)<0$.
For all $q\in\mathbf{N}$, let
$A^{\flat}_{q}=\{r\in\mathbf{N}:r\geq q\}$
and let
$A_{q}=A\cup A^{\flat}_{q}=A\cup\{r\in\mathbf{N}:r\geq q\}.$
The sequence $(A_{q})_{q=1}^{\infty}$ is decreasing and $A=\bigcap_{q=1}^{\infty}A_{q}$.
If $q\geq 1$ and $q\notin A$, then $q\in A_{q}$ but $q\notin A_{q+1}$ and so $A_{q}\neq A_{q+1}$. Because AA does not contain all sufficiently large integers, the sequence $(A_{q})_{q=1}^{\infty}$ is asymptotically strictly decreasing.
For all $h\in\mathbf{N}$ and $q\in\mathbf{N}$, we have
$hA\subseteq hA_{q}\subseteq hA\cup\{r\in\mathbf{N}:r\geq q+(h-1)a^{*}\}$
and
$\bigcap_{q=1}^{\infty}\{r\in\mathbf{N}:r\geq q+(h-1)a^{*}\}=\emptyset.$
Therefore,
$hA\subseteq\bigcap_{q=1}^{\infty}hA_{q}\subseteq hA\cup\bigcap_{q=1}^{\infty}\{r\in\mathbf{N}:r\geq q+(h-1)a^{*}\}=hA.$
This completes the proof. ∎
Problem 2.
Let AA be an infinite set of integers that is not a basis of order $h$ for $\mathbf{Z}$. Does there exist a strictly decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ and $hA=\bigcap_{q=1}^{\infty}hA_{q}$?
Problem 3.
For what sets AA of integers does there exist a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ and $hA\neq\bigcap_{q=1}^{\infty}hA_{q}$ for all $h\in\mathbf{N}$?
Problem 4.
For what sets AA of integers does there exist an asymptotically decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $hA=\bigcap_{q=1}^{\infty}hA_{q}$ for all $h\in\mathbf{N}$?
3. Sumsets in locally compact groups
Theorem 4.
Let AA be a nonempty subset of a locally compact abelian group $G$. Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of closed compact subsets of $G$ such that $A=\bigcap_{q=1}^{\infty}A_{q}$. Then
$hA=\bigcap_{q=1}^{\infty}hA_{q}$
for every positive integer $h$.
Proof.
For all $h\in\mathbf{N}$ and $q\in\mathbf{N}$, the sumset $hA_{q}$ and the sumset intersection $\bigcap_{q=1}^{\infty}hA_{q}$ are nonempty closed compact sets. Let $x\in\bigcap_{q=1}^{\infty}hA_{q}$. We must show that $x\in hA$.
For all $q\in\mathbf{N}$, we have $x\in hA_{q}$ and so there is an $h$-tuple
$(a_{1,q},\ldots,a_{h,q})\in A_{q}^{h}\subseteq A_{1}^{h}$
such that
$x=a_{1,q}+\cdots+a_{h,q}.$
Consider the sequences $(a_{i,q})_{q=1}^{\infty}$ for $i=1,\ldots,h$. Because the set $A_{1}$ is compact and $a_{i,q}\in A_{1}$ for all $i=1,\ldots,h$ and $q\in\mathbf{N}$, there is an infinite strictly increasing sequence of positive integers $(q_{j})_{j=1}^{\infty}$ such that the $h$ subsequences $(a_{i,q_{j}})_{j=1}^{\infty}$ converge:
$\lim_{j\rightarrow\infty}a_{i,q_{j}}=x_{i}$
for all $i=1,\ldots,h$. Then
$x$ $=\lim_{j\rightarrow\infty}\left(a_{1,q_{j}}+\cdots+a_{h,q_{j}}\right)$
$=\left(\lim_{j\rightarrow\infty}a_{1,q_{j}}\right)+\cdots+\left(\lim_{j\rightarrow\infty}a_{h,q_{j}}\right)$
$=x_{1}+\cdots+x_{h}.$
Suppose that $x_{i}\notin A$ for some ii. Let $q_{0}=0$. Construct the sequence $\left(x_{i,q}\right)_{q=1}^{\infty}$ as follows;
$x_{i,q}=a_{i,q_{j}}\qquad\text{for $q\in[q_{j-1}+1,q_{j}]$.}$
Because $(A_{q})_{q=1}^{\infty}$ is a decreasing sequence, we have $x_{i,q}\in A_{q}$ for all $q\in\mathbf{N}$ and
$\lim_{i\rightarrow\infty}x_{i,q}=x_{i}.$
The group $G$ is locally compact. Because AA is closed and $x_{i}\notin A$, there is a compact neighborhood $K$ of $x_{i}$ that is disjoint from AA such that $x_{i,q}\in K$ for all sufficiently large $q$. Then $x_{i,q}\in K\cap A_{q}$ for all sufficiently large $q$. It follows that, for all $Q\in\mathbf{N}$, there exists $q_{1}\geq Q$ with $x_{i,q_{1}}\in K\cap A_{q_{1}}\subseteq K\cap A_{q}$ for all $q\in[1,Q]$, and so
$x_{i,q_{1}}\in\bigcap_{q=1}^{Q}\left(K\cap A_{q}\right)\neq\emptyset.$
Because the sets $K\cap A_{q}$ are compact, the finite intersection property implies
$\emptyset\neq\bigcap_{q=1}^{\infty}\left(K\cap A_{q}\right)=K\cap\left(\bigcap_{q=1}^{\infty}A_{q}\right)=K\cap A=\emptyset$
which is absurd. Therefore, $x_{i}\in A$ for all $i=1,\ldots,h$ and $h\in hA$. This completes the proof. ∎
Theorem 5.
Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of nonempty compact sets in $\mathbf{R}^{n}$ and let $A=\bigcap_{q=1}^{\infty}A_{q}$. For every positive integer $h$,
$hA=\bigcap_{q=1}^{\infty}hA_{q}.$
Proof.
This follows from Theorem 4 with $G=\mathbf{R}^{n}$. ∎
Theorem 6.
Let $G$ be a locally compact abelian group with Haar measure $\mu$. For $h\in\mathbf{N}$, let $\theta_{h}\geq 0$ and let $(\theta_{h,q})_{q=1}^{\infty}$ be a sequence of nonnegative real numbers such that $\lim_{q\rightarrow\infty}\theta_{h,q}=\theta_{h}$. Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of compact subsets of $G$ such that $\mu(hA_{q})=\theta_{h,q}$ for all $q\in\mathbf{N}$. Let $A=\bigcap_{q=1}^{\infty}A_{q}$. Then $\mu(hA)=\theta_{h}$.
Proof.
For all $h\in\mathbf{N}$, the sequence $(hA_{q})_{q=1}^{\infty}$ is a decreasing sequence of compact subsets of $G$. By Theorem 4,
$hA=\bigcap_{q=1}^{\infty}hA_{q}$
and so
$\mu(hA)=\mu\left(\bigcap_{q=1}^{\infty}hA_{q}\right)=\lim_{q\rightarrow\infty}\mu(hA_{q})=\lim_{q\rightarrow\infty}\mu(\theta_{h,q})=\theta_{h}.$
This completes the proof. ∎
4. Semigroup problems
Problem 5.
Let $(A_{q})_{q=1}^{\infty}$ be a strictly decreasing sequence of sets in an additive abelian semigroup and let $A=\bigcap_{q=1}^{\infty}A_{q}$. Describe the set $\{h\in\mathbf{N}:hA=\bigcap_{q=1}^{\infty}hA_{q}\}$.
Problem 6.
Let $(A_{q})_{q=1}^{\infty}$ be a strictly decreasing sequence of sets in an additive abelian semigroup and let $A=\bigcap_{q=1}^{\infty}A_{q}$.
When does $hA=\bigcap_{q=1}^{\infty}hA_{q}$ imply $(h+1)A=\bigcap_{q=1}^{\infty}(h+1)A_{q}$?
When does $hA=\bigcap_{q=1}^{\infty}hA_{q}$ imply $(h-1)A=\bigcap_{q=1}^{\infty}(h-1)A_{q}$?