### v2

# Intersections of sumsets in additive number theory

Abstract.
Let $A$ be a subset of an additive abelian semigroup and let $hA$ be the $h$-fold sumset of $A$. The following question is considered: Let $(A_{q})_{q=1}^{\infty}$ be a strictly decreasing sequence of sets in the semigroup and let $A=\bigcap_{q=1}^{\infty}A_{q}$. When does one have
$hA=\bigcap_{q=1}^{\infty}hA_{q}$
for some or all $h\geq 2$?
Key words and phrases:
Sumset, Haar measure of sumsets, additive number theory, combinatorial number theory
2000 Mathematics Subject Classification:
11B13, 11B05, 11B75, 11P70, 22D99
Supported in part by PSC-CUNY Research Award Program grant 66197-00 54.
1. The sumset intersection problem
Let $\mathbf{N}$, $\mathbf{N}_{0}$, and $\mathbf{Z}$ denote, respectively, the sets of positive integers, nonnegative integers, and integers.
Let $S$ be an additive abelian semigroup and let $A$ be a subset of $S$. For every positive integer $h$, the $h$-fold sumset of $A$ is the set of all sums of $h$ not necessarily distinct elements of $A$, that is,
$hA=\{a_{1}+\cdots+a_{h}:a_{i}\in A\text{ for all }i=1,\ldots,h\}.$
For all $x\in S$, we denote by $r_{A,h}(x)$ the number of representations of $x$ as a sum of $h$ elements of $A$, that is,
$r_{A,h}(x)=\operatorname{\text{card}}\left\{(a_{1},\ldots,a_{h})\in A^{h}:a_{1}+\cdots+a_{h}=x\right\}.$
The subset $A$ is a basis of order $h$ for $S$ if $hA=S$ (equivalently, if $r_{A,h}(x)\geq 1$ for all $x\in S$), and a basis for $S$ if $A$ is a basis of order $h$ for $S$ for some positive integer $h$. The subset $A$ is a nonbasis of order $h$ for $S$ if $A$ is not a basis of order $h$ for $S$, that is, if $hA\neq S$. The set $A$ is a nonbasis for $S$ if $A$ is a nonbasis of order $h$ for $S$ for every positive integer $h$.
The sequence of sets $(A_{q})_{q=1}^{\infty}$ is decreasing if $A_{q}\supseteq A_{q+1}$ for all $q\in\mathbf{N}$, and strictly decreasing if $A_{q}\neq A_{q+1}$ for all $q\in\mathbf{N}$. The sequence $(A_{q})_{q=1}^{\infty}$ asymptotically strictly decreasing if the sequence is not eventually constant, that is, if $A_{q}\neq A_{q+1}$ for infinitely many $q$. Every asymptotically strictly decreasing sequence contains a strictly decreasing subsequence.
Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of sets in the semigroup $S$ and let
$A=\bigcap_{q=1}^{\infty}A_{q}.$
For all $q\in\mathbf{N}$, we have $A\subseteq A_{q}$ and so $hA\subseteq hA_{q}$. It follows that
(1) $hA=h\left(\bigcap_{q=1}^{\infty}A_{q}\right)\subseteq\bigcap_{q=1}^{\infty}hA_{q}.$
This inclusion can be proper. For example, in the semigroup $\mathbf{Z}$ of integers, for every positive integer $q$, we define
$A^{\sharp}_{q}=\{r\in\mathbf{Z}:|r|\geq q\}.$
Let $h\geq 2$. For every nonnegative integer $n$, we have
$n=(n+q)+(h-2)q+(-(h-1)q)\in hA^{\sharp}_{q}$
and
$n=(-n-q)+(h-2)(-q)+((h-1)q)\in hA^{\sharp}_{q}$
and so $hA^{\sharp}_{q}=\mathbf{Z}$.
Let $A$ be any finite set of integers. The sumset $hA$ is finite for all $h\in\mathbf{N}$ and so $hA\neq\mathbf{Z}$. The asymptotically strictly decreasing sequence of sets $(A_{q})_{q=1}^{\infty}$ defined by
$A_{q}=A\cup A^{\sharp}_{q}=A\cup\{r\in\mathbf{Z}:|r|\geq q\}$
satisfies $A=\bigcap_{q=1}^{\infty}A_{q}$. Because $hA_{q}=\mathbf{Z}$ for all $q\in\mathbf{N}$ , we have
$hA\neq\mathbf{Z}=\bigcap_{q=1}^{\infty}hA_{q}.$
We want to know when equality replaces inclusion in relation (1).
Problem 1.
Describe, classify, understand the strictly decreasing sequences $(A_{q})_{q=1}^{\infty}$ of subsets of an additive abelian semigroup such that
(2) $h\left(\bigcap_{q=1}^{\infty}A_{q}\right)=\bigcap_{q=1}^{\infty}hA_{q}.$
For related work, see [1][11].
2. Intersections of sets of integers
Theorem 1.
Let $A$ be a set of nonnegative integers. Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of sets of nonnegative integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$. For every positive integer $h$,
$hA=\bigcap_{q=1}^{\infty}hA_{q}.$
Proof.
We must show that for all $h\geq 2$, if $n\in\bigcap_{q=1}^{\infty}hA_{q}$, then $n\in hA$.
The number $r_{\mathbf{N}_{0},h}(n)$ of representations of $n$ as a sum of $h$ nonnegative integers is finite. Indeed, the set $\mathcal{R}(n)$ of $h$-tuples $(b_{1},\ldots,b_{h})\in\mathbf{N}_{0}^{h}$ such that
$n=b_{1}+\cdots+b_{h}$
satisfies $r_{\mathbf{N}_{0},h}(n)=|\mathcal{R}(n)|=\binom{n+h-1}{h-1}$. For all $q\in\mathbf{N}$, we have $n\in hA_{q}$ and so there is an $h$-tuple $(a_{1,q},\ldots,a_{h,q})\in A_{q}^{h}\cap\mathcal{R}(n)$ such that
$n=a_{1,q}+\cdots+a_{h,q}.$
Because $\mathcal{R}(n)$ is finite, there is some $h$-tuple $(b_{1},\ldots,b_{h})\in\mathcal{R}(n)$ such that
$(b_{1},\ldots,b_{h})=(a_{1,q},\ldots,a_{h,q})$
for infinitely many $q$ and so, for infinitely many $q$, we have $b_{i}=a_{i,q}\in A_{q}$ for all $i=1,\ldots,h$. Because the sequence $(A_{q})_{q=1}^{\infty}$ is decreasing, we have $b_{i}\in A_{q}$ for all $i=1,\ldots,h$ and for all $q$. It follows that $b_{i}\in\bigcap_{q=1}^{\infty}A_{q}=A$ and $n=b_{1}+\cdots+b_{h}\in hA$. This completes the proof. ∎
Theorem 2.
Let $S$ be an additive abelian semigroup such that $r_{S,h}(x)$ is finite for all $h\in\mathbf{N}$ and $x\in S$. Let $A$ be a subset of $S$ and let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of subsets of $S$ such that $A=\bigcap_{q=1}^{\infty}A_{q}$. For every positive integer $h$,
$hA=\bigcap_{q=1}^{\infty}hA_{q}.$
Proof.
It suffices to observe that the proof of Theorem 1 uses only the finiteness of the representation function $R_{\mathbf{N}_{0},h}(n)$ for all $h$ and $n$. This result follows immediately. ∎
Theorem 3.
Let $S$ be an additive abelian semigroup and let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of subsets of $S$ such that $A=\bigcap_{q=1}^{\infty}A_{q}$. If $A$ is a basis of order $h_{0}$ for $S$, then
$hA=\bigcap_{q=1}^{\infty}hA_{q}$
for all $h\geq h_{0}$.
Proof.
For all $h\geq h_{0}$ and for all $q\in\mathbf{N}$, we have $hA=hA_{q}=S$. and so
$hA=\mathbf{Z}=\bigcap_{q=1}^{\infty}hA_{q}.$
This completes the proof. ∎
Theorem 4.
Let $A$ be a set of integers. If $A$ is a nonbasis of order $h_{0}$ for $\mathbf{Z}$, then there exists a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ but
$h_{0}A\neq\bigcap_{q=1}^{\infty}hA_{q}.$
If $A$ is a nonbasis for $\mathbf{Z}$, then there exists a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ but
$hA\neq\bigcap_{q=1}^{\infty}hA_{q}$
for all $h\in\mathbf{N}$.
Proof.
For all $q\in\mathbf{N}$, let
$A_{q}=A\cup A_{q}^{\sharp}=A\cup\{r\in\mathbf{Z}:|r|\geq q\}.$
Then $(A_{q})_{q=1}^{\infty}$ is a decreasing sequence of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$. Moreover, $h_{0}A_{q}=\mathbf{Z}$ for all $q\in\mathbf{N}$. If $A$ is a nonbasis of order $h_{0}$ for $\mathbf{Z}$, then
$h_{0}A\neq\mathbf{Z}=\bigcap_{q=1}^{\infty}h_{0}A_{q}.$
If $A$ is a nonbasis for $\mathbf{Z}$, then for all $h\in\mathbf{N}$ we have
$hA\neq\mathbf{Z}=\bigcap_{q=1}^{\infty}hA_{q}.$
This completes the proof. ∎
Theorem 4 extends to the group of lattice points $\mathbf{Z}^{d}$ and to a more general class of semigroups. For the group $\mathbf{Z}^{d}$, we let
$A_{q}=\left\{(n_{1},\ldots,n_{d})\in\mathbf{Z}^{d}:|n_{i}|\geq q\text{ for all }q\in\mathbf{N}\right\}.$
Theorem 5.
Let $S$ be an additive abelian semigroup such that $S$ contains a strictly decreasing sequence of sets $\left(A_{q}^{\sharp}\right)_{q=1}^{\infty}$ such that $\bigcap_{q=1}^{\infty}A_{q}^{\sharp}=\emptyset$ and $hA_{q}^{\sharp}=S$ for all $h\geq 2$ and $q\geq 1$.
If $A$ is a nonbasis of order $h_{0}$ for $S$, then there exists a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ but
$h_{0}A\neq\bigcap_{q=1}^{\infty}hA_{q}.$
If $A$ is a nonbasis for $S$, then there exists a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ but
$hA\neq\bigcap_{q=1}^{\infty}hA_{q}$
for all $h\in S$.
Proof.
If $A$ is a nonbasis of order $h_{0}$ for $S$, then $h_{0}A\neq S$. For all $q\in\mathbf{N}$, let
$A_{q}=A\cup A_{q}^{\sharp}.$
Then $A=\bigcap_{q=1}^{\infty}A_{q}$ and $h_{0}A_{q}=S$ for all $q\in\mathbf{N}$. It follows that
$h_{0}A\neq S=\bigcap_{q=1}^{\infty}h_{0}A_{q}.$
If $A$ is a nonbasis for $S$, then for all $h\in\mathbf{N}$ we have $hA\neq S$ and
$hA\neq S=\bigcap_{q=1}^{\infty}hA_{q}$
for all $h\in\mathbf{N}$. This completes the proof. ∎
Theorem 6.
Let $A$ be an infinite set of integers that is bounded below and that does not contain all sufficiently large integers. There exists an asymptotically strictly decreasing sequence $(A_{q})_{q=1}^{\infty}$ such that $hA=\bigcap_{q=1}^{\infty}hA_{q}$ for all $h\in\mathbf{N}$.
Proof.
By Theorem 1, we can assume that $a^{*}=\min(A)<0$.
For all $q\in\mathbf{N}$, let
$A^{\flat}_{q}=\{r\in\mathbf{N}:r\geq q\}$
and let
$A_{q}=A\cup A^{\flat}_{q}=A\cup\{r\in\mathbf{N}:r\geq q\}.$
The sequence $(A_{q})_{q=1}^{\infty}$ is decreasing and $A=\bigcap_{q=1}^{\infty}A_{q}$.
If $q\geq 1$ and $q\notin A$, then $q\in A_{q}$ but $q\notin A_{q+1}$ and so $A_{q}\neq A_{q+1}$. Because $A$ does not contain all sufficiently large integers, the sequence $(A_{q})_{q=1}^{\infty}$ is asymptotically strictly decreasing.
For all $h\in\mathbf{N}$ and $q\in\mathbf{N}$, we have
$hA\subseteq hA_{q}\subseteq hA\cup\{r\in\mathbf{N}:r\geq q+(h-1)a^{*}\}$
and
$\bigcap_{q=1}^{\infty}\{r\in\mathbf{N}:r\geq q+(h-1)a^{*}\}=\emptyset.$
Therefore,
$hA\subseteq\bigcap_{q=1}^{\infty}hA_{q}\subseteq hA\cup\bigcap_{q=1}^{\infty}\{r\in\mathbf{N}:r\geq q+(h-1)a^{*}\}=hA.$
This completes the proof. ∎
Problem 2.
Let $A$ be an infinite set of integers that is a nonbasis of order $h$ for $\mathbf{Z}$. Does there exist a strictly decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ and $hA=\bigcap_{q=1}^{\infty}hA_{q}$?
Problem 3.
For what sets $A$ of integers does there exist a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ and $hA\neq\bigcap_{q=1}^{\infty}hA_{q}$ for all $h\in\mathbf{N}$?
Problem 4.
For what sets $A$ of integers does there exist an asymptotically decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $hA=\bigcap_{q=1}^{\infty}hA_{q}$ for all $h\in\mathbf{N}$?
3. Sumsets in locally compact groups
Theorem 7.
Let $A$ be a nonempty subset of a locally compact abelian group $G$. Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of closed compact subsets of $G$ such that $A=\bigcap_{q=1}^{\infty}A_{q}$. Then
$hA=\bigcap_{q=1}^{\infty}hA_{q}$
for every positive integer $h$.
Proof.
For all $h\in\mathbf{N}$ and $q\in\mathbf{N}$, the sumset $hA_{q}$ and the sumset intersection $\bigcap_{q=1}^{\infty}hA_{q}$ are nonempty closed compact sets. Let $x\in\bigcap_{q=1}^{\infty}hA_{q}$. We must show that $x\in hA$.
For all $q\in\mathbf{N}$, we have $x\in hA_{q}$ and so there is an $h$-tuple
$(a_{1,q},\ldots,a_{h,q})\in A_{q}^{h}\subseteq A_{1}^{h}$
such that
$x=a_{1,q}+\cdots+a_{h,q}.$
Consider the sequences $(a_{i,q})_{q=1}^{\infty}$ for $i=1,\ldots,h$. Because the set $A_{1}$ is compact and $a_{i,q}\in A_{1}$ for all $i=1,\ldots,h$ and $q\in\mathbf{N}$, there is an infinite strictly increasing sequence of positive integers $(q_{j})_{j=1}^{\infty}$ such that the $h$ subsequences $(a_{i,q_{j}})_{j=1}^{\infty}$ converge:
$\lim_{j\rightarrow\infty}a_{i,q_{j}}=x_{i}$
for all $i=1,\ldots,h$. Then
$x$ $=\lim_{j\rightarrow\infty}\left(a_{1,q_{j}}+\cdots+a_{h,q_{j}}\right)$
$=\left(\lim_{j\rightarrow\infty}a_{1,q_{j}}\right)+\cdots+\left(\lim_{j\rightarrow\infty}a_{h,q_{j}}\right)$
$=x_{1}+\cdots+x_{h}.$
Suppose that $x_{i}\notin A$ for some $i$. Let $q_{0}=0$. Construct the sequence $\left(x_{i,q}\right)_{q=1}^{\infty}$ as follows;
$x_{i,q}=a_{i,q_{j}}\qquad\text{for $q\in[q_{j-1}+1,q_{j}]$.}$
Because $(A_{q})_{q=1}^{\infty}$ is a decreasing sequence, we have $x_{i,q}\in A_{q}$ for all $q\in\mathbf{N}$ and
$\lim_{i\rightarrow\infty}x_{i,q}=x_{i}.$
The group $G$ is locally compact. Because $A$ is closed and $x_{i}\notin A$, there is a compact neighborhood $K$ of $x_{i}$ that is disjoint from $A$ such that $x_{i,q}\in K$ for all sufficiently large $q$. Then $x_{i,q}\in K\cap A_{q}$ for all sufficiently large $q$. It follows that, for all $Q\in\mathbf{N}$, there exists $q_{1}\geq Q$ with $x_{i,q_{1}}\in K\cap A_{q_{1}}\subseteq K\cap A_{q}$ for all $q\in[1,Q]$, and so
$x_{i,q_{1}}\in\bigcap_{q=1}^{Q}\left(K\cap A_{q}\right)\neq\emptyset.$
Because the sets $K\cap A_{q}$ are compact, the finite intersection property implies
$\emptyset\neq\bigcap_{q=1}^{\infty}\left(K\cap A_{q}\right)=K\cap\left(\bigcap_{q=1}^{\infty}A_{q}\right)=K\cap A=\emptyset$
which is absurd. Therefore, $x_{i}\in A$ for all $i=1,\ldots,h$ and $h\in hA$. This completes the proof. ∎
Theorem 8.
Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of nonempty compact sets in $\mathbf{R}^{n}$ and let $A=\bigcap_{q=1}^{\infty}A_{q}$. For every positive integer $h$,
$hA=\bigcap_{q=1}^{\infty}hA_{q}.$
Proof.
This follows from Theorem 7 with $G=\mathbf{R}^{n}$. ∎
Theorem 9.
Let $G$ be a locally compact abelian group with Haar measure $\mu$. For $h\in\mathbf{N}$, let $\theta_{h}\geq 0$ and let $(\theta_{h,q})_{q=1}^{\infty}$ be a sequence of nonnegative real numbers such that $\lim_{q\rightarrow\infty}\theta_{h,q}=\theta_{h}$. Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of compact subsets of $G$ such that $\mu(hA_{q})=\theta_{h,q}$ for all $q\in\mathbf{N}$. Let $A=\bigcap_{q=1}^{\infty}A_{q}$. Then $\mu(hA)=\theta_{h}$.
Proof.
For all $h\in\mathbf{N}$, the sequence $(hA_{q})_{q=1}^{\infty}$ is a decreasing sequence of compact subsets of $G$. By Theorem 7,
$hA=\bigcap_{q=1}^{\infty}hA_{q}$
and so
$\mu(hA)=\mu\left(\bigcap_{q=1}^{\infty}hA_{q}\right)=\lim_{q\rightarrow\infty}\mu(hA_{q})=\lim_{q\rightarrow\infty}\mu(\theta_{h,q})=\theta_{h}.$
This completes the proof. ∎
4. Semigroup problems
Problem 5.
Let $(A_{q})_{q=1}^{\infty}$ be a strictly decreasing sequence of sets in an additive abelian semigroup and let $A=\bigcap_{q=1}^{\infty}A_{q}$. Describe the set $\{h\in\mathbf{N}:hA=\bigcap_{q=1}^{\infty}hA_{q}\}$.
Problem 6.
Let $(A_{q})_{q=1}^{\infty}$ be a strictly decreasing sequence of sets in an additive abelian semigroup and let $A=\bigcap_{q=1}^{\infty}A_{q}$.
When does $hA=\bigcap_{q=1}^{\infty}hA_{q}$ imply $(h+1)A=\bigcap_{q=1}^{\infty}(h+1)A_{q}$?
When does $hA=\bigcap_{q=1}^{\infty}hA_{q}$ imply $(h-1)A=\bigcap_{q=1}^{\infty}(h-1)A_{q}$?
I thank Diego Marques for the observation that led to Theorem 2.