### v3

# Intersections of sumsets in additive number theory

Abstract.
Let $A$ be a subset of an additive abelian semigroup $S$ and let $hA$ be the $h$-fold sumset of $A$. The following question is considered: Let $(A_{q})_{q=1}^{\infty}$ be a strictly decreasing sequence of sets in $S$ and let $A=\bigcap_{q=1}^{\infty}A_{q}$. When does one have
$hA=\bigcap_{q=1}^{\infty}hA_{q}$
for some or all $h\geq 2$?
Key words and phrases:
Sumsets in semigroups, Haar measure of sumsets, additive number theory, combinatorial number theory
2000 Mathematics Subject Classification:
11B13, 11B05, 11B75, 11P70, 22D99
Supported in part by PSC-CUNY Research Award Program grant 66197-00 54.
1. The sumset intersection problem
Let $\mathbf{N}$, $\mathbf{N}_{0}$, and $\mathbf{Z}$ denote, respectively, the sets of positive integers, nonnegative integers, and integers.
Let $S$ be an additive abelian semigroup and let $A$ be a subset of $S$. For every positive integer $h$, the $h$-fold sumset of $A$ is the set of all sums of $h$ not necessarily distinct elements of $A$, that is,
$hA=\{a_{1}+\cdots+a_{h}:a_{i}\in A\text{ for all }i=1,\ldots,h\}.$
For all $x\in S$, we denote by $r_{A,h}(x)$ the number of representations of $x$ as a sum of $h$ elements of $A$, that is,
$r_{A,h}(x)=\operatorname{\text{card}}\left\{(a_{1},\ldots,a_{h})\in A^{h}:a_{1}+\cdots+a_{h}=x\right\}.$
The subset $A$ is a basis of order $h$ for $S$ if $hA=S$, or, equivalently, if $r_{A,h}(x)\geq 1$ for all $x\in S$. The subset $A$ is a basis for $S$ if $A$ is a basis of order $h$ for $S$ for some positive integer $h$. The subset $A$ is a nonbasis of order $h$ for $S$ if $A$ is not a basis of order $h$ for $S$, that is, if $hA\neq S$. The set $A$ is a nonbasis for $S$ if $hA\neq S$ for every positive integer $h$.
The sequence of sets $(A_{q})_{q=1}^{\infty}$ is decreasing if $A_{q}\supseteq A_{q+1}$ for all $q\in\mathbf{N}$. The decreasing sequence $(A_{q})_{q=1}^{\infty}$ is strictly decreasing if $A_{q}\neq A_{q+1}$ for all $q\in\mathbf{N}$ and asymptotically strictly decreasing if $A_{q}\neq A_{q+1}$ for infinitely many $q$. A decreasing sequence is asymptotically strictly decreasing if and only if it is not eventually constant. Every asymptotically strictly decreasing sequence contains a strictly decreasing subsequence.
Let $(A_{q})_{q=1}^{\infty}$ be an asymptotically strictly decreasing sequence of sets in the semigroup $S$ and let
$A=\bigcap_{q=1}^{\infty}A_{q}.$
For all $q\in\mathbf{N}$, we have $A\subseteq A_{q}$ and so $hA\subseteq hA_{q}$. It follows that
(1) $hA=h\left(\bigcap_{q=1}^{\infty}A_{q}\right)\subseteq\bigcap_{q=1}^{\infty}hA_{q}.$
This inclusion can be proper. For example, in the semigroup $\mathbf{Z}$ of integers, for every positive integer $q$, we define
(2) $A^{\sharp}_{q}=\{r\in\mathbf{Z}:|r|\geq q\}.$
Let $h\geq 2$. For every nonnegative integer $n$, we have $\pm(n+q)\in A^{\sharp}_{q}$ and $\pm(h-1)q\in A^{\sharp}_{q}$. It follows that
$n=(n+q)+(h-2)q+(-(h-1)q)\in hA^{\sharp}_{q}$
and
$n=-(n+q)+(h-2)(-q)+((h-1)q)\in hA^{\sharp}_{q}$
and so $hA^{\sharp}_{q}=\mathbf{Z}$.
Let $A$ be any finite set of integers. The sumset $hA$ is finite for all $h\in\mathbf{N}$ and so $hA\neq\mathbf{Z}$. The asymptotically strictly decreasing sequence of sets $(A_{q})_{q=1}^{\infty}$ defined by
$A_{q}=A\cup A^{\sharp}_{q}=A\cup\{r\in\mathbf{Z}:|r|\geq q\}$
satisfies $A=\bigcap_{q=1}^{\infty}A_{q}$. Because $hA_{q}=\mathbf{Z}$ for all $q\in\mathbf{N}$ , we have
$hA\neq\mathbf{Z}=\bigcap_{q=1}^{\infty}hA_{q}.$
Problem 1.
We want to know when equality replaces inclusion in relation (1). The basic problem is to describe, classify, understand the strictly decreasing sequences $(A_{q})_{q=1}^{\infty}$ of subsets of an additive abelian semigroup such that
(3) $h\left(\bigcap_{q=1}^{\infty}A_{q}\right)=\bigcap_{q=1}^{\infty}hA_{q}.$
For related work, see [2][13].
2. Intersections of sumsets in discrete groups and semigroups
Theorem 1.
Let $h\in\mathbf{N}$. Let $S$ be an additive abelian semigroup such that $r_{S,h}(x)$ is finite for all $x\in S$. Let $A$ be a subset of $S$ and let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of subsets of $S$ such that $A=\bigcap_{q=1}^{\infty}A_{q}$. Then
$hA=\bigcap_{q=1}^{\infty}hA_{q}.$
Proof.
For all $x\in S$, the set $\mathcal{R}_{S,h}(x)$ of $h$-tuples $(b_{1},\ldots,b_{h})\in S^{h}$ such that
$x=b_{1}+\cdots+b_{h}$
is finite with cardinality $|\mathcal{R}_{S,h}(x)|=r_{S,h}(x)$.
Let $x\in\bigcap_{q=1}^{\infty}hA_{q}$. We must show that $x\in hA$. For all $q\in\mathbf{N}$, we have $x\in hA_{q}$ and so there is an $h$-tuple $(a_{1,q},\ldots,a_{h,q})\in A_{q}^{h}\cap\mathcal{R}_{S,h}(x)$ such that
$x=a_{1,q}+\cdots+a_{h,q}.$
Because $\mathcal{R}_{S,h}(x)$ is a finite set, there is some $h$-tuple $(b_{1},\ldots,b_{h})\in\mathcal{R}_{S,h}(x)$ such that
$(b_{1},\ldots,b_{h})=(a_{1,q},\ldots,a_{h,q})$
for infinitely many $q$ and so, for infinitely many $q$, we have $b_{i}=a_{i,q}\in A_{q}$ for all $i=1,\ldots,h$. Because the sequence $(A_{q})_{q=1}^{\infty}$ is decreasing, we have $b_{i}\in A_{q}$ for all $i=1,\ldots,h$ and $q\in\mathbf{N}$. It follows that $b_{i}\in\bigcap_{q=1}^{\infty}A_{q}=A$ and $x=b_{1}+\cdots+b_{h}\in hA$. This completes the proof. ∎
Let $\mathbf{Z}^{d}$ be the additive abelian group of $d$-dimensional lattice points in $\mathbf{R}^{d}$ and let $\mathbf{N}_{0}^{d}$ be the additive abelian semigroup of $d$-dimensional nonnegative lattice points.
Theorem 2.
Let $d\geq 1$ and let $A$ be a subset of $\mathbf{N}_{0}^{d}$. Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of subsets of $\mathbf{N}_{0}^{d}$ such that $A=\bigcap_{q=1}^{\infty}A_{q}$. For every positive integer $h$,
$hA=\bigcap_{q=1}^{\infty}hA_{q}.$
Proof.
For all $n\in\mathbf{N}_{0}$, the number $r_{\mathbf{N}_{0},h}(n)$ of representations of $n$ as a sum of $h$ nonnegative integers is finite. Indeed, $r_{\mathbf{N}_{0},h}(n)=\binom{n+h-1}{h-1}$. For all $\mathbf{n}=(n_{1},\ldots,n_{d})\in\mathbf{N}_{0}^{d}$, we have
$r_{\mathbf{N}_{0}^{d},h}(\mathbf{n})=\prod_{i=1}^{d}r_{\mathbf{N}_{o},h}(n_{i})<\infty.$
The result follows from Theorem 1. ∎
Theorem 3.
Let $S$ be an additive abelian semigroup and let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of subsets of $S$ such that $A=\bigcap_{q=1}^{\infty}A_{q}$. If $A$ is a basis of order $h_{0}$ for $S$, then
$h_{0}A=\bigcap_{q=1}^{\infty}h_{0}A_{q}$
If the subgroup $S$ contains an additive identity and if $A$ is a basis of order $h_{0}$ for $S$, then
$hA=\bigcap_{q=1}^{\infty}hA_{q}$
for all $h\geq h_{0}$.
Proof.
For all $q\in\mathbf{N}$, we have $A\subseteq A_{q}$ and so $S=h_{0}A\subseteq h_{0}A_{q}\subseteq S$. It follows that $h_{0}A_{q}=S$ and
$h_{0}A=S=\bigcap_{q=1}^{\infty}hA_{q}.$
Let $h\geq h_{0}$. If $S$ contains an additive identity, then $h_{0}A\subseteq hA$. If $A$ is a basis of order $h_{0}$ for $S$, then $S=h_{0}A\subseteq hA\subseteq S$ and so $A$ is a basis of order $h$ for $S$. This completes the proof. ∎
Note that the additive semigroup $\mathbf{N}$ of positive integers does not contain an additive identity. The set $A=\mathbf{N}$ is a basis of order 1 for $\mathbf{N}$ but not a basis of order $h$ for any $h\geq 2$.
Theorem 4.
Let $d\in\mathbf{N}$ and let $G$ be any subgroup of $\mathbf{R}^{d}$ that contains $\mathbf{Z}^{d}$. Let $A$ be a subset of $G$. If $A$ is a nonbasis of order $h_{0}$ for $G$, then there exists a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of subsets of $G$ such that $A=\bigcap_{q=1}^{\infty}A_{q}$ and
$h_{0}A\neq\bigcap_{q=1}^{\infty}h_{0}A_{q}.$
If $A$ is a nonbasis for $G$, then there exists a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of subsets of $G$ such that $A=\bigcap_{q=1}^{\infty}A_{q}$ and
$hA\neq\bigcap_{q=1}^{\infty}hA_{q}$
for all $h\geq 2$.
Proof.
For all $\mathbf{x}=(x_{1},\ldots,x_{d})\in\mathbf{R}^{d}$, define $\|\mathbf{x}\|=\max(|x_{i}|:i=1,\ldots,d)$. For all $q\in\mathbf{N}$, let
$A_{q,d}^{\sharp}=\{\mathbf{x}\in G:\|\mathbf{x}\|\geq q\}.$
Note that $\{\mathbf{x}\in\mathbf{Z}^{d}:\|\mathbf{x}\|\geq q\}\subseteq A_{q,d}^{\sharp}$.
Let $\mathbf{x}=(x_{1},\ldots,x_{d})\in G$. If $x_{1}\geq 0$, then
$\mathbf{y}=(x_{1}+(h-1)q,x_{2},\ldots,x_{d})\in A_{q}^{\sharp}$
and
$\mathbf{z}=(-q,0,\ldots,0)\in A_{q}^{\sharp}$
and so
$\mathbf{x}$ $=(x_{1}+(h-1)q,x_{2},\ldots,x_{d})+(h-1)(-q,0,\ldots,0)$
$=\mathbf{y}+(h-1)\mathbf{z}\in hA_{q,d}^{\sharp}.$
If $x_{1}\leq 0$, then
$(x_{1}-(h-1)q,x_{2},\ldots,x_{d})\in A_{q}^{\sharp}$
and
$(q,0,\ldots,0)\in A_{q}^{\sharp}$
and so
$\mathbf{x}$ $=(x_{1}-(h-1)q,x_{2},\ldots,x_{d})+(h-1)(q,0,\ldots,0)$
$=\mathbf{y}+(h-1)\mathbf{z}\in hA_{q,d}^{\sharp}.$
Thus, $hA_{q,d}^{\sharp}=G$ for all $q\in\mathbf{N}$ and $h\geq 2$.
For all $q\in\mathbf{N}$, let
$A_{q}=A\cup A_{q,d}^{\sharp}.$
Then $hA_{q}=hA_{q,d}^{\sharp}=G$ for all $q\in\mathbf{N}$ and $h\geq 2$. The sequence $(A_{q})_{q=1}^{\infty}$ is a decreasing sequence of sets such that $A=\bigcap_{q=1}^{\infty}A_{q}$. If $A$ is a nonbasis of order $h_{0}$ for $G$, then
$h_{0}A\neq G=\bigcap_{q=1}^{\infty}h_{0}A_{q}.$
If $A$ is a nonbasis for $G$, then
$hA\neq G=\bigcap_{q=1}^{\infty}hA_{q}$
for all $h\geq 2$. This completes the proof. ∎
Theorem 5.
Let $A$ be an infinite set of integers that is bounded below and that does not contain all sufficiently large integers. There exists an asymptotically strictly decreasing sequence $(A_{q})_{q=1}^{\infty}$ such that $hA=\bigcap_{q=1}^{\infty}hA_{q}$ for all $h\in\mathbf{N}$.
Proof.
Note that a set of integers that is bounded below cannot be a basis for $\mathbf{Z}$. By Theorem 2 with $d=1$, we can assume that $a^{*}=\min(A)<0$.
For all $q\in\mathbf{N}$, let
$A^{\flat}_{q}=\{r\in\mathbf{N}:r\geq q\}$
and let
$A_{q}=A\cup A^{\flat}_{q}=A\cup\{r\in\mathbf{N}:r\geq q\}.$
The sequence $(A_{q})_{q=1}^{\infty}$ is decreasing and $A=\bigcap_{q=1}^{\infty}A_{q}$. If $q\in\mathbf{N}$ and $q\notin A$, then $q\in A_{q}$ but $q\notin A_{q+1}$ and so $A_{q}\neq A_{q+1}$. Because $A$ does not contain all sufficiently large integers, the sequence $(A_{q})_{q=1}^{\infty}$ is asymptotically strictly decreasing.
For all $h\in\mathbf{N}$ and $q\in\mathbf{N}$, we have
$hA\subseteq hA_{q}\subseteq hA\cup\{r\in\mathbf{N}:r\geq q+(h-1)a^{*}\}$
and
$\bigcap_{q=1}^{\infty}\{r\in\mathbf{N}:r\geq q+(h-1)a^{*}\}=\emptyset.$
Therefore,
$hA\subseteq\bigcap_{q=1}^{\infty}hA_{q}\subseteq hA\cup\bigcap_{q=1}^{\infty}\{r\in\mathbf{N}:r\geq q+(h-1)a^{*}\}=hA.$
This completes the proof. ∎
Let $S$ be an additive abelian semigroup. A subset $A$ of $S$ is a maximal nonbasis of order $h$ for $S$ if $A$ is not a basis of order $h$ for $S$ but every proper superset of $A$ is a basis of order $h$ for $S$. For example, for every integer $h\geq 2$, the set $h\ast\mathbf{Z}=\{h\ell:\ell\in\mathbf{Z}\}$ of multiples of $h$ is a maximal nonbasis of order $h$ for $\mathbf{Z}$ if and only if $h$ is prime. Maximal nonbases were introduced by Nathanson [4] and the first nontrivial examples were constructed by Nathanson [4] and Erdős and Nathanson [1].
Theorem 6.
Let $A$ be a subset of the infinite abelian semigroup $S$. If $A$ is a maximal nonbasis of order $h$ for $S$ and if $(A_{q})_{q=1}^{\infty}$ is an asymptotically strictly decreasing sequence of subsets of $S$ such that $A=\bigcap_{q=1}^{\infty}A_{q}$, then $hA\neq\bigcap_{q=1}^{\infty}hA_{q}$.
Proof.
Let $A$ be a maximal nonbasis of order $h$ for $S$ and let $(A_{q})_{q=1}^{\infty}$ be a strictly decreasing sequence such that $A=\bigcap_{q=1}^{\infty}A_{q}$. For all $q$, the set $A$ is a proper subset of $A_{q}$ and so $hA_{q}=S$. It follows that $hA\neq S=\bigcap_{q=1}^{\infty}hA_{q}$. This completes the proof. ∎
Problem 2.
Let $A$ be a subset of the infinite abelian semigroup $S$ such that $A$ is a nonbasis of order $h$ but not a maximal nonbasis of order $h$ for $S$. Does there exist an asymptotically strictly decreasing sequence $(A_{q})_{q=1}^{\infty}$ of subsets of $S$ such that $A=\bigcap_{q=1}^{\infty}A_{q}$ and $hA=\bigcap_{q=1}^{\infty}hA_{q}$? In particular, consider this problem for $S=\mathbf{Z}$.
3. Intersections of sumsets in locally compact groups
Theorem 7.
Let $A$ be a nonempty subset of a locally compact abelian group $G$. Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of compact subsets of $G$ such that $A=\bigcap_{q=1}^{\infty}A_{q}$. Then
$hA=\bigcap_{q=1}^{\infty}hA_{q}$
for every positive integer $h$.
Proof.
For all $h\in\mathbf{N}$ and $q\in\mathbf{N}$, the sumset $hA_{q}$ and the sumset intersection $\bigcap_{q=1}^{\infty}hA_{q}$ are nonempty closed compact sets. Let $x\in\bigcap_{q=1}^{\infty}hA_{q}$. We must show that $x\in hA$.
For all $q\in\mathbf{N}$, we have $x\in hA_{q}$ and so there is an $h$-tuple
$(a_{1,q},\ldots,a_{h,q})\in A_{q}^{h}\subseteq A_{1}^{h}$
such that
$x=a_{1,q}+\cdots+a_{h,q}.$
Consider the sequences $(a_{i,q})_{q=1}^{\infty}$ for $i=1,\ldots,h$. Because the set $A_{1}$ is compact and $a_{i,q}\in A_{1}$ for all $i=1,\ldots,h$ and $q\in\mathbf{N}$, there is an infinite strictly increasing sequence of positive integers $(q_{j})_{j=1}^{\infty}$ such that the $h$ subsequences $(a_{i,q_{j}})_{j=1}^{\infty}$ converge:
$\lim_{j\rightarrow\infty}a_{i,q_{j}}=x_{i}$
for all $i=1,\ldots,h$. Then
$x$ $=\lim_{j\rightarrow\infty}\left(a_{1,q_{j}}+\cdots+a_{h,q_{j}}\right)$
$=\left(\lim_{j\rightarrow\infty}a_{1,q_{j}}\right)+\cdots+\left(\lim_{j\rightarrow\infty}a_{h,q_{j}}\right)$
$=x_{1}+\cdots+x_{h}.$
Suppose that $x_{i}\notin A$ for some $i$. Let $q_{0}=0$. Construct the sequence $\left(x_{i,q}\right)_{q=1}^{\infty}$ as follows;
$x_{i,q}=a_{i,q_{j}}\qquad\text{for $q\in[q_{j-1}+1,q_{j}]$.}$
Because $(A_{q})_{q=1}^{\infty}$ is a decreasing sequence, we have $x_{i,q}\in A_{q}$ for all $q\in\mathbf{N}$ and
$\lim_{i\rightarrow\infty}x_{i,q}=x_{i}.$
The group $G$ is locally compact. Because $A=\bigcap_{q=1}^{\infty}A_{q}$ is closed and $x_{i}\notin A$, there is a compact neighborhood $K$ of $x_{i}$ that is disjoint from $A$ such that $x_{i,q}\in K$ for all sufficiently large $q$. Then $x_{i,q}\in K\cap A_{q}$ for all sufficiently large $q$. It follows that, for all $Q\in\mathbf{N}$, there exists $q_{1}\geq Q$ with $x_{i,q_{1}}\in K\cap A_{q_{1}}\subseteq K\cap A_{q}$ for all $q\in[1,Q]$, and so
$x_{i,q_{1}}\in\bigcap_{q=1}^{Q}\left(K\cap A_{q}\right)\neq\emptyset.$
Because the sets $K\cap A_{q}$ are compact, the finite intersection property implies
$\emptyset\neq\bigcap_{q=1}^{\infty}\left(K\cap A_{q}\right)=K\cap\left(\bigcap_{q=1}^{\infty}A_{q}\right)=K\cap A=\emptyset$
which is absurd. Therefore, $x_{i}\in A$ for all $i=1,\ldots,h$ and $x\in hA$. This completes the proof. ∎
Theorem 8.
Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of nonempty compact sets in $\mathbf{R}^{d}$ such that $A=\bigcap_{q=1}^{\infty}A_{q}\neq\emptyset$. For every positive integer $h$,
$hA=\bigcap_{q=1}^{\infty}hA_{q}.$
Proof.
This follows from Theorem 7 with $G=\mathbf{R}^{d}$. ∎
Theorem 9.
Let $G$ be a locally compact abelian group with Haar measure $\mu$. For $h\in\mathbf{N}$, let $\theta_{h}\geq 0$ and let $(\theta_{h,q})_{q=1}^{\infty}$ be a decreasing sequence of nonnegative real numbers such that $\lim_{q\rightarrow\infty}\theta_{h,q}=\theta_{h}$. Let $(A_{q})_{q=1}^{\infty}$ be a decreasing sequence of compact subsets of $G$ such that $\mu(hA_{q})=\theta_{h,q}$ for all $q\in\mathbf{N}$. Let $A=\bigcap_{q=1}^{\infty}A_{q}$. Then $\mu(hA)=\theta_{h}$.
Proof.
For all $h\in\mathbf{N}$, the sequence $(hA_{q})_{q=1}^{\infty}$ is a decreasing sequence of compact subsets of $G$. By Theorem 7,
$hA=\bigcap_{q=1}^{\infty}hA_{q}$
and so
$\mu(hA)=\mu\left(\bigcap_{q=1}^{\infty}hA_{q}\right)=\lim_{q\rightarrow\infty}\mu(hA_{q})=\lim_{q\rightarrow\infty}\theta_{h,q}=\theta_{h}.$
This completes the proof. ∎
4. Additional problems
Problem 3.
Let $(A_{q})_{q=1}^{\infty}$ be a strictly decreasing sequence of sets in an additive abelian semigroup and let $A=\bigcap_{q=1}^{\infty}A_{q}$. Describe the set $\{h\in\mathbf{N}:hA=\bigcap_{q=1}^{\infty}hA_{q}\}$. What sets of integers arise in this way?
Problem 4.
Let $(A_{q})_{q=1}^{\infty}$ be a strictly decreasing sequence of sets in an additive abelian semigroup and let $A=\bigcap_{q=1}^{\infty}A_{q}$.
When does $hA=\bigcap_{q=1}^{\infty}hA_{q}$ imply $(h+1)A=\bigcap_{q=1}^{\infty}(h+1)A_{q}$?
When does $hA=\bigcap_{q=1}^{\infty}hA_{q}$ imply $(h-1)A=\bigcap_{q=1}^{\infty}(h-1)A_{q}$?
Problem 5.
For what sets $A$ of integers does there exist a decreasing sequence $(A_{q})_{q=1}^{\infty}$ of sets of integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$ and $hA\neq\bigcap_{q=1}^{\infty}hA_{q}$ for all $h\in\mathbf{N}$?
Acknowledgement. I thank Diego Marques for helpful comments that suggested Theorem 6.